OpenStudy (anonymous):

Let f(x) = 2-x and g(x) = 1/x. Perform each function operation and then find the domain of the result. g/f(x)

hero (hero):

I thought I helped you with this already.

OpenStudy (anonymous):

would it be 1/2x-x^2

OpenStudy (anonymous):

equate denominator=0 and omit that x values for domain

hero (hero):

Yes, and by the way: (General Rule) \[\frac{a}{b} \div \frac{b}{a} = \frac{a}{b} \times \frac{a}{b}\]. In other words, flip the second fraction

hero (hero):

Yes, and make sure that the denominator doesn't equal zero

hero (hero):

\[2x - x^2 \ne 0\]

OpenStudy (anonymous):

domain={R/x not equal to 0 and 2}

OpenStudy (anonymous):

(-\[(-\infty,0]U[0,2]U[2,\infty)\]

OpenStudy (anonymous):

thangiee's answer completly confused me.

OpenStudy (anonymous):

\[(1)/(2x-x^2)\]

hero (hero):

\[x(2-x) \ne 0\]\[x \ne 0 \space \text{and} \space x \ne -2\]

hero (hero):

^ That refers to the domain @trolololcat

hero (hero):

Remember, we're saying that the not equal sign means that x cannot equal zero and x cannot equal negative 2 because that would make the denominator equal zero, which is a "no-no" in math because we cannot divide by zero....ever! :D

OpenStudy (anonymous):

1/0 is undefined

hero (hero):

So basically, what we mean is that when you post your solution to (g/f)(x), include what the domain cannot equal as part of your answer.

OpenStudy (anonymous):

To find the domain you always set the denominator to =0 and solve?

hero (hero):

Yes

OpenStudy (anonymous):

yes but should omit that x values

hero (hero):

And that would help you find out what x cannot equal. I prefer to write it with the not equal sign

hero (hero):

Which I demonstrated earlier

OpenStudy (anonymous):

ok so when you're dividing you multiply and when you multiple you divide? for this case

hero (hero):

Actually, to re-correct myself: \[x \ne 0 \space \text{and} \space x \ne 2\]

hero (hero):

Don't confuse yourself. I told you to go and review what happens when you have a case like \[\frac{1}{2} \div 3\] For some reason, you think you're doing something different than basic math when you learned how to multiply by the reciprocal when dividing fractions.

OpenStudy (anonymous):

wait, & for \[(x-1)^2/x \] x=1? or it's just left like that fration

OpenStudy (anonymous):

I'll look over it. I never really liked fractions, i always tried turning anything into decimals lol

hero (hero):

The only thing you have to worry about with regard to the domain and fractions is to make sure that x or any expression in the denominator cannot equal zero.

OpenStudy (anonymous):

Can you help me with one more? @Hero f(x)=2x^2+x-3 g(x)=x-1 (f*g)(x) So it would be set up like (2x^2+x-3)(x-1) correct?

OpenStudy (anonymous):

@Hero

hero (hero):

Yes, correct. And the domain is all real numbers because no fraction or square root

hero (hero):

Don't forget to multiply and simplify

OpenStudy (anonymous):

2x^3+x^2+2x-3

OpenStudy (anonymous):

2x^3-x^2+2x-3 I mean

hero (hero):

That's probably right. Anyway, I have to go. I want to get a few more episodes of 24 in before I hit the hay.

hero (hero):